Our teacher gave us a worksheet and there's some questions I don't understand. One is this... The sum of an integer and three more than three times the next consecutive integer is 30. Find the integers.
And we are using the 5 step method and I just can't get all the steps right, except number one. Please help.

Answer by mananth(16949)

Integer = x
next integer = x+1
3 more than 3 times the next integer 3(x+1) +3
Sum of both = 30
x+3(x+1)+3=30
x+3x+3+3 =30
4x+6 =30
-6 -6
4x=24
/4
x= 6
6 and 7 are the integers